# Error-Minimizing Measurements in Postselected One-Shot Symmetric Quantum State Discrimination and Acceptance as a Performance Metric

### v1

@@ 3. For case $\underline{\mathcal{C}\mathrm{3}}$,
Here $c$ and $c_{r}$ are as given in the Table I for an arbitrary error-minimizing measurement.
@@
Using Theorem C.1 from Appendix C, we have
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathrm{Tr}\left(\sigma^{1/2}\psi_{\max}\sigma^{1/2}\sigma^{-1/2}\rho\sigma^{-1/2}\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}\mathrm{Tr}\left(\sigma^{1/2}\psi_{\max}\sigma^{1/2}\mathrm{T}^{\max}\sigma^{-1/2}\rho\sigma^{-1/2}\right)$
=$\stackrel{{\scriptstyle(d)}}{{=}}R_{\max}(\rho,\sigma)\mathrm{Tr}\left(\sigma^{1/2}\psi_{\max}\sigma^{1/2}\mathrm{T}^{\max}\right)$
=$\stackrel{{\scriptstyle(e)}}{{=}}R_{\max}(\rho,\sigma)\mathrm{Tr}\left(\sigma^{1/2}\psi_{\max}\sigma^{1/2}\right).$
Here, $(a)$ uses the condition that $\Pi_{\sigma}=\Pi_{\rho}$. $(b)$ is obtained from $\Pi_{\sigma}=\sigma^{-1/2}\sigma^{1/2}$ and using cyclic property of trace. $(c)$ and $(e)$ uses that $\sigma^{1/2}\psi_{\max}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$. $(d)$ follows from the fact that $\mathrm{T}^{\max}$ is projection onto the subspace corresponding to maximum eigenvalue of $\sigma^{-1/2}\rho\sigma^{-1/2}$ and so, we have $\mathrm{T}^{\max}\sigma^{-1/2}\rho\sigma^{-1/2}=R_{\max}(\rho,\sigma)\mathrm{T}^{\max}$. Thus, we obtain
@@
In a similar fashion, we can show that
@@
We will use these results from (25),(26), and (27) directly for $\Gamma=\Lambda_{\rho}$ and $\Gamma=\Lambda_{\sigma}$. The set of measurements for which error is minimized are given differently for the three cases, so we derive expression of acceptance for the three cases one-by-one as given below.
@@ 1. $\underline{\mathcal{C}\mathrm{1}}$
    $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ are parameterized as $\Lambda_{\rho}=c\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)},\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho}$ and corresponding acceptance is
@@ 2. $\underline{\mathcal{C}\mathrm{2}}$
    $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ are parameterized as $\Lambda_{\sigma}=c\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)},\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma}$ and corresponding acceptance is
@@ 3. $\underline{\mathcal{C}\mathrm{3}}$
    $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ are parameterized as $\Lambda_{\rho}=cc_{r}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)},\Lambda_{\sigma}=c(1-c_{r})\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)},\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$ and corresponding acceptance is given as
@@
Barring the case $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)=R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, the expression for acceptance only depends on $c$ linearly. In turn, $c$ is upper bounded by a function of $\psi_{\max},\psi_{\min}$ and $c_{r}$. So, for maximizing acceptance, we only need to focus on maximizing $c$ over the set of all error minimizing measurements.
For these identified POVMs that achieve the minimum postselected symmetric error, the maximum acceptance is given in the following theorem.
Theorem III.4.
Under the constraint $e(\Lambda)=e_{s}(\rho,\sigma,p),\Lambda\in\mathcal{M}$, the maximum acceptance for the states $\rho$ and $\sigma$ is given below.
@@
Here, for any $\mathrm{Tr}\left(\Pi_{1}\Pi_{2}\right)=0$, we have
$\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)\stackrel{{\scriptstyle\Delta}}{{=}}\{\max c:c(r\psi_{1}+(1-r)\psi_{2})\leq\sigma\text{ for some }\psi_{1}\in\mathcal{S}(\Pi_{1}),\psi_{2}\in\mathcal{S}(\Pi_{2})\}.$
More detailed properties of $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)$ are given in Appendix E, which would help finding closed form expression in specific cases.
@@
We derive for the three cases separately.
@@ 1. $\underline{\mathcal{C}\mathrm{1}}$
    From Lemma III.3, $A^{s}_{\rho}=R_{\max}(\rho,\sigma)\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c\qquad\text{ and }\qquad A^{s}_{\sigma}=\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c.$ Now $\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c$ is obtained from Lemma D.2 in Appendix D as stated below
@@
    Substituting it in the expression of $A_{\rho}^{s}$ and $A_{\sigma}^{s}$, we get the stated results.
@@ 2. $\underline{\mathcal{C}\mathrm{2}}$
    From Lemma III.3, $A^{s}_{\rho}=R_{\min}(\rho,\sigma)\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c\qquad\text{ and }\qquad A^{s}_{\sigma}=\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c.$ Now $\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c$ is obtained from Lemma D.2 in Appendix D as stated below
@@
    Substituting it in the expression of $A_{\rho}^{s}$ and $A_{\sigma}^{s}$, we get the stated results.
@@ 3. $\underline{\mathcal{C}\mathrm{3}}$
    From Lemma III.3, $A_{\rho}(\Lambda)=cc_{r}R_{\max}(\rho,\sigma)+c(1-c_{r})R_{\min}(\rho,\sigma)\text{ and }A_{\sigma}(\Lambda)=c.$ Note that these functions have to be maximized with respect to $c$ and $c_{r}$. Taking first and maximizing $c$ we get
@@
    Here the last step is obtained from Lemma D.3 in Appendix D. Now maximizing $A_{\sigma}(\Lambda)$, we obtain
@@
    Here the last step is obtained from Lemma D.3 in Appendix D.
@@
Note that for the case $\mathcal{C}1$, the expressions for acceptance for $\rho$ and $\sigma$, both are linear in the parameter $c$ in Theorem III.3, so the parameters that maximize acceptance for both states $\rho$ and $\sigma$ is the same. Thus, same measurement maximizes both the acceptances. This also holds for the case $\mathcal{C}2$. In contrast, in the case $\mathcal{C}3$, we obtained acceptances as different functions of $c$ and $c_{r}$. So, the measurements which maximize acceptance for the state $\rho,i.e.A_{\rho}(\Lambda)$, won’t be maximizing acceptance for the state $\sigma,i.e.A_{\sigma}(\Lambda)$ and the measurement which maximize the acceptance for the state $\sigma,i.e.A_{\sigma}(\Lambda)$ won’t be maximizing acceptance for the state $\rho,i.e.A_{\rho}(\Lambda)$.
@@
The maximum acceptance can be obtained by taking the measurement described by the POVM $\Lambda=\{\mathrm{P}^{\max},0,\mathrm{I}-\mathrm{P}^{\max}\}$ and $\Lambda=\{0,\mathrm{P}^{\min},\mathrm{I}-\mathrm{P}^{\min}\}$ in $\mathcal{C}1$ and $\mathcal{C}2$ respectively. In $\mathcal{C}3$, finding the measurement, which achieves the maximum acceptance for the state $\rho$ (or $\sigma$) is involved. First step is finding $c_{r}$ such that $A_{\rho}^{s}$ (or $A_{\sigma}^{s}$) is obtained and placing corresponding optimal $\psi_{\max}$ and $\psi_{\min}$ and $c=\Upsilon_{\mathrm{T}^{\max},\mathrm{T}^{\min}}(\sigma,c_{r})$, maximum acceptance achieving measurement can be obtained. This requires solving the optimization step and finding the solution.
This completes deriving the expression for maximum acceptance over the set of measurements which achieves minimum postselected symmetric error. In the next section, we maximize the acceptance for the pair of states $\rho$ and $\sigma$ such that $\Pi_{\rho}\neq\Pi_{\sigma}$.
IV The possible states $\rho$ and $\sigma$ do not have the same support i.e. $\Pi_{\rho}\neq\Pi_{\sigma}$
@@
Recall from (9) that the minimum postselected symmetric error $e_{s}(\rho,\sigma,p)=\left(\Xi\left(p_{\rho}\rho,p_{\sigma}\sigma\right)+1\right)^{-1}$. If $\Pi_{\rho}\neq\Pi_{\sigma}$, from the definition given in (8), it turns out that $\Xi\left(p_{\rho}\rho,p_{\sigma}\sigma\right)=\infty$, hence $e_{s}(\rho,\sigma,p)=0$. In this section, we begin with finding the condition on $\Lambda$ to obtain $e(\Lambda)=e_{s}(\rho,\sigma,p)=0$. The following theorem states the set of all possible error-minimizing measurements as a union of three sets, with each corresponding to one of the three conditions needed to ensure that the error vanishes. An arbitrary measurement must belong to one of three sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ or $\mathcal{E}_{s}^{3}(\rho,\sigma)$ defined in the Theorem below to ensure that $e(\Lambda)=0$.
Theorem IV.1.
For any measurement $\Lambda\in\mathcal{M}$, $e(\Lambda)=0$ if and only if
$\Lambda\in\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{1}_{s}(\rho,\sigma)\cup\mathcal{E}^{2}_{s}(\rho,\sigma)\cup\mathcal{E}^{3}_{s}(\rho,\sigma),$ (28)
@@
where $\mathcal{E}^{1}_{s}(\rho,\sigma)\stackrel{{\scriptstyle\Delta}}{{=}}\{\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\}$, $\mathcal{E}^{2}_{s}(\rho,\sigma)\stackrel{{\scriptstyle\Delta}}{{=}}\{\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\}$, $\mathcal{E}^{3}_{s}(\rho,\sigma)\stackrel{{\scriptstyle\Delta}}{{=}}\{\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\}$.
@@
The proof is given in Appendix F. ∎
@@
The three sets described in the Theorem IV.1 have a specific structure of outcomes, which is described below (also see Fig. 5).
@@
    The first set $\mathcal{E}_{s}^{1}(\rho,\sigma)$ consists of all the measurements for which only $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)$ i.e. probability of detecting $\rho$ when the given state is $\rho$ is non-vanishing and the probabilities $\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)$, $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)$ and $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$ vanish. So, for any measurement that belongs to this set, if the given input state is $\rho$, the measurement either declares in favor of $\rho$ or rejects both the hypotheses. However, if the given state is $\sigma$, it always rejects both the hypotheses. Fig. 5(a) illustrates this point.
@@
    Similarly, the second set i.e. $\mathcal{E}_{s}^{2}(\rho,\sigma)$ consists of all the measurements for which only $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$ i.e. probability of detecting $\sigma$ when the given state is $\sigma$ is not vanishing and the probabilities $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)$, $\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)$, and $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)$ vanish. So, for any measurement that belongs to this set, if the given input state is $\sigma$, the measurement either declares in favor of $\sigma$ or rejects both the hypotheses. However, if the given state is $\rho$, it always rejects both the hypotheses. Fig. 5(b) illustrates this point.
@@
Note that for any measurement for which error vanishes, $\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)=0$ or equivalently $\mathrm{Tr}(\Lambda_{\rho}\sigma)=0$ and $\mathrm{Tr}(\Lambda_{\sigma}\rho)=0$ and so both type-1 and type-2 error has to vanish. This follows from the fact that numerator of $e(\Lambda)$ has to vanishes to ensure that error vanishes. So, it either declares the correct outcome or reject both the hypotheses but never declares the wrong outcome. Further note that, one of $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)$ or $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$ i.e. at least one of the probability of correct decision has to be non-zero to ensure that denominator in $e(\Lambda)$ remains non-zero.
@@
Note that the set and respective conditions depend only the subspace spanned by eigenvectors of $\rho$ and $\sigma$ and not on the prior probabilities $p_{\rho}$ and $p_{\sigma}$. So, error-minimizing measurements are independent of the prior probabilities when $\Pi_{\rho}\neq\Pi_{\sigma}$.

### v2

@@ 3. For case $\underline{\mathcal{C}\mathrm{3}}$,
Here, $c$ and $c_{r}$ depend on $\Lambda$ and are as given in the Table III.
@@
Using Theorem B.1 from Appendix B, we have
@@
$\stackrel{{\scriptstyle(b)}}{{=}}\mathrm{Tr}\left(\sigma^{1/2}\psi_{\max}\sigma^{1/2}\sigma^{-1/2}\rho\sigma^{-1/2}\right)$
$\stackrel{{\scriptstyle(c)}}{{=}}\mathrm{Tr}\left(\sigma^{1/2}\psi_{\max}\sigma^{1/2}\mathrm{T}^{\max}\sigma^{-1/2}\rho\sigma^{-1/2}\right)$
$\stackrel{{\scriptstyle(d)}}{{=}}R_{\max}(\rho,\sigma)\mathrm{Tr}\left(\sigma^{1/2}\psi_{\max}\sigma^{1/2}\mathrm{T}^{\max}\right)$
$\stackrel{{\scriptstyle(e)}}{{=}}R_{\max}(\rho,\sigma)\mathrm{Tr}\left(\sigma^{1/2}\psi_{\max}\sigma^{1/2}\right).$
Here, $(a)$ is due to $\Pi_{\sigma}=\Pi_{\rho}$. $(b)$ is obtained from $\Pi_{\sigma}=\sigma^{-1/2}\sigma^{1/2}$ and using cyclic property of trace. $(c)$ and $(e)$ uses that $\sigma^{1/2}\psi_{\max}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$. $(d)$ follows from the fact that $\mathrm{T}^{\max}$ is the projection onto the subspace corresponding to the maximum eigenvalue of $\sigma^{-1/2}\rho\sigma^{-1/2}$ and so, we have $\mathrm{T}^{\max}\sigma^{-1/2}\rho\sigma^{-1/2}=R_{\max}(\rho,\sigma)\mathrm{T}^{\max}$. Thus, we obtain
@@
Similarly, we can show that
@@
We will use these results from (25), (26), and (27) directly for $\Gamma=\Lambda_{\rho}$ and $\Gamma=\Lambda_{\sigma}$. Now, for three cases as given in Theorem 1, acceptance is given as follows.
@@ 1. $\underline{\mathcal{C}\mathrm{1}}$
    Recall that $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ are parameterized as $\Lambda_{\rho}=c\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)},\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho}$. Hence, corresponding acceptance is
@@ 2. $\underline{\mathcal{C}\mathrm{2}}$
    Recall that $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ are parameterized as $\Lambda_{\sigma}=c\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)},\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma}$. Hence, corresponding acceptance is
@@ 3. $\underline{\mathcal{C}\mathrm{3}}$
    Recall that $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ are parameterized as $\Lambda_{\rho}=cc_{r}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)},\Lambda_{\sigma}=c(1-c_{r})\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)},\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$. Hence, corresponding acceptance is given as
@@
Barring the case $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)=R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, the expression for acceptance depends only on $c$ linearly. In turn, $c$ is upper bounded by a function of $\psi_{\max},\psi_{\min}$ and $c_{r}$. So, to maximize acceptance, we only need to focus on maximizing $c$ for all error-minimizing measurements.
We now give the maximum value of acceptance that an error-minimizing measurement can have, in the following theorem.
Theorem 4.
For an error-minimizing measurement $\Lambda$, (i.e., $e(\Lambda)=e_{s}(\rho,\sigma,p),\Lambda\in\mathcal{M}$), the maximum value of acceptance for the states $\rho$ and $\sigma$ is given below.
@@
Here, $\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)$ is defined as
$\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)\stackrel{{\scriptstyle\Delta}}{{=}}\{\max c:c(r\psi_{1}+(1-r)\psi_{2})\leq\sigma\text{ for some }\psi_{1}\in\mathcal{S}(\Pi_{1}),\psi_{2}\in\mathcal{S}(\Pi_{2}),\text{ with }\mathrm{Tr}\left(\Pi_{1}\Pi_{2}\right)=0\}.$
More detailed properties of $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)$ are given in Appendix D, which can help find closed-form expressions of maximum acceptance in specific cases.
@@
Let us take the three cases separately.
@@ 1. $\underline{\mathcal{C}\mathrm{1}}$
    From Lemma 3, $A^{s}_{\rho}=R_{\max}(\rho,\sigma)\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c\qquad\text{ and }\qquad A^{s}_{\sigma}=\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c.$ Now, from Lemma C.1 in Appendix C, we have
@@
    Substituting it in the expression of $A_{\rho}^{s}$ and $A_{\sigma}^{s}$, as given in Theorem 3, we get the result stated in the theorem.
@@ 2. $\underline{\mathcal{C}\mathrm{2}}$
    From Lemma 3, $A^{s}_{\rho}=R_{\min}(\rho,\sigma)\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c\qquad\text{ and }\qquad A^{s}_{\sigma}=\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c.$ Now, from Lemma C.1 in Appendix C, we have
@@
    Substituting it in the expression of $A_{\rho}^{s}$ and $A_{\sigma}^{s}$, as given in Theorem 3, we get the result stated in the theorem.
@@ 3. $\underline{\mathcal{C}\mathrm{3}}$
    From Theorem 3, $A_{\rho}(\Lambda)=cc_{r}R_{\max}(\rho,\sigma)+c(1-c_{r})R_{\min}(\rho,\sigma)\text{ and }A_{\sigma}(\Lambda)=c.$ Note that these functions have to be maximized with respect to $c$ and $c_{r}$. Taking $c$ common and maximizing it first, we get
@@
    Here, the last step is obtained from Lemma C.2 in Appendix C. Now maximizing $A_{\sigma}(\Lambda)$, we obtain
@@
    Here, the last step is due to Lemma C.2 in Appendix C.
@@
Note that for any desired value of acceptance, error-minimizing measurements can be designed by choosing the appropriate value of constant $c$ and $c_{r}$, provided that the desired acceptance remains less than the maximum obtained in Theorem 4.
@@
Note that for the case $\mathcal{C}1$, the expressions of acceptance for $\rho$ and $\sigma$, both are linear in the parameter $c$ in Theorem 3, so the parameter-values that maximize acceptance for both states $\rho$ and $\sigma$ is the same. Thus, the same measurements maximize both acceptances. This observation also holds for the case $\mathcal{C}2$. In contrast, in the case $\mathcal{C}3$, we obtained acceptances as different functions of $c$ and $c_{r}$. So, the measurements that maximize acceptance for the state $\rho,i.e.,A_{\rho}(\Lambda)$, will not be maximizing acceptance for the state $\sigma,i.e.,A_{\sigma}(\Lambda)$. Similarly, the measurement that maximizes the acceptance for the state $\sigma,i.e.,A_{\sigma}(\Lambda)$ will not maximize acceptance for the state $\rho,i.e.,A_{\rho}(\Lambda)$.
Remark 8.
@@
The maximum acceptance for the two state can simultaneously be obtained by taking the measurement described by the POVM $\Lambda=\{\mathrm{P}^{\max},0,\mathrm{I}-\mathrm{P}^{\max}\}$ and $\Lambda=\{0,\mathrm{P}^{\min},\mathrm{I}-\mathrm{P}^{\min}\}$ in $\mathcal{C}1$ and $\mathcal{C}2$ respectively. In $\mathcal{C}3$, finding the measurement that achieves the maximum acceptance for the state $\rho$ (or $\sigma$) is involved. First step is finding $c_{r}$ such that $A_{\rho}^{s}$ (or $A_{\sigma}^{s}$) is obtained and placing corresponding optimal $\psi_{\max}$ and $\psi_{\min}$ and $c=\Upsilon_{\mathrm{T}^{\max},\mathrm{T}^{\min}}(\sigma,c_{r})$, the maximum acceptance achieving measurement can be obtained by solving the optimization step. In $\mathcal{C}3$, the measurements that maximize acceptance for state $\rho$ and $\sigma$ are different.
Now, having completed the analysis for the pair of states with the same support, in the next section, we maximize the acceptance for the pair of states $\rho$ and $\sigma$ such that $\Pi_{\rho}\neq\Pi_{\sigma}$.
IV Case: $\rho$ and $\sigma$ do not have the same support i.e. $\Pi_{\rho}\neq\Pi_{\sigma}$
@@
Recall from (9) that the minimum postselected symmetric error $e_{s}(\rho,\sigma,p)=\left(\Xi\left(p_{\rho}\rho,p_{\sigma}\sigma\right)+1\right)^{-1}$. If $\Pi_{\rho}\neq\Pi_{\sigma}$, from the definition given in (8), it turns out that $\Xi\left(p_{\rho}\rho,p_{\sigma}\sigma\right)=\infty$, hence $e_{s}(\rho,\sigma,p)=0$. In this section, we begin with finding the condition on $\Lambda$ to obtain $e(\Lambda)=e_{s}(\rho,\sigma,p)=0$. The following theorem states the set of all possible error-minimizing measurements as a union of three sets, each corresponding to one of the three conditions that ensures that the error vanishes. An arbitrary measurement must belong to one of three sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ or $\mathcal{E}_{s}^{3}(\rho,\sigma)$ defined in the Theorem below to ensure that $e(\Lambda)=0$. A parameterization of measurements in these sets is described later in Lemma 1, 2 and 3 to allow construction of an arbitrary measurement from each of these sets.
Theorem 5.
For any measurement $\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}\in\mathcal{M}$, $e(\Lambda)=0$ if and only if
[…]
@@
$\mathcal{E}^{1}_{s}(\rho,\sigma)$ $\stackrel{{\scriptstyle\Delta}}{{=}}\{\Lambda:\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\},$
$\mathcal{E}^{2}_{s}(\rho,\sigma)$ $\stackrel{{\scriptstyle\Delta}}{{=}}\{\Lambda:\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\},$
$\mathcal{E}^{3}_{s}(\rho,\sigma)$ $\stackrel{{\scriptstyle\Delta}}{{=}}\{\Lambda:\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\}.$
@@
The proof is given in Appendix E. ∎
@@
The three sets described in the Theorem 5 have a specific structure of outcomes, which is described below (also see Fig. 5).
@@
    The first set $\mathcal{E}_{s}^{1}(\rho,\sigma)$ consists of all the measurements for which only $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)$ i.e. the probability of detecting $\rho$ when the given state is $\rho$ is non-zero and the probabilities $\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)$, $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)$ and $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$ vanish. So, for any measurement that belongs to this set, if the given input state is $\rho$, the measurement either declares in favor of $\rho$ or doesn’t make any decision (i.e. reject the test). However, if the given state is $\sigma$, it always rejects both the hypotheses. Fig. 5(a) illustrates this point.
@@
    Similarly, the second set i.e. $\mathcal{E}_{s}^{2}(\rho,\sigma)$ consists of all the measurements for which only $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$ i.e. the probability of detecting $\sigma$ when the given state is $\sigma$ is non-zero and the probabilities $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)$, $\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)$, and $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)$ vanish. So, for any measurement that belongs to this set, if the given input state is $\sigma$, the measurement either declares in favor of $\sigma$ or reject the test. However, if the given state is $\rho$, it always rejects both the hypotheses. Fig. 5(b) illustrates this point.
@@
Note that for any measurement for which error vanishes, $\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)=0$, or equivalently $\mathrm{Tr}(\Lambda_{\rho}\sigma)=0$ and $\mathrm{Tr}(\Lambda_{\sigma}\rho)=0$. Thus, both type-1 and type-2 error must vanish. It follows from the fact that the numerator of $e(\Lambda)$ has to vanish to ensure that the error vanishes. So, such a measurement either declares the correct outcome or rejects both hypotheses but never declares the wrong outcome. Quantum state discrimination with such requirements is called unambiguous state discrimination (USD). Measurements that lead to zero probability of wrong outcome are called USD measurements. Such discrimination has been studied quite a lot in the past. On taking the number of states to be discriminated as $2$, USD measurements obtained in [9, 13, 37, 4, 26] belong to the set given in Theorem 5. Further, note that one of $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)$ or $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$, i.e., at least one of the probability of correct decision has to be non-zero to ensure that the denominator in $e(\Lambda)$ remains non-zero.
@@
Note that the set and respective conditions depend only on the subspace spanned by eigenvectors of $\rho$ and $\sigma$ and not on the prior probabilities $p_{\rho}$ and $p_{\sigma}$. So, error-minimizing measurements are independent of the prior probabilities when $\Pi_{\rho}\neq\Pi_{\sigma}$.